OpenStudy (anonymous):

Find the minimum value of x where sin(2x+45)=cos(30-x), using the factor theorem.

OpenStudy (anonymous):

Factor theorem? Isn't this sum and difference formulae?

OpenStudy (anonymous):

Yes. It's under the same name in Malaysia.

OpenStudy (anonymous):

ima write it out on paper, too much typing >.> lol

myininaya (myininaya):

\[cos(\frac{\pi}{2}-(2x+45))=\sin(2x+45)\] so \[\cos(\frac{\pi}{2}-(2x+45))=\cos(30-x)\] \[\frac{\pi}{2}-(2x+45)=30-x\]

OpenStudy (anonymous):

>.> too pro!

myininaya (myininaya):

lol

myininaya (myininaya):

i have to work today :(

myininaya (myininaya):

sort of i have to sit through a meeting that is forever

OpenStudy (anonymous):

eewww meetings >.<

OpenStudy (anonymous):

im gonna tell the high school I work at im putting in my two weeks.

myininaya (myininaya):

poor high school they need and love joe

OpenStudy (akshay_budhkar):

myininaya meeting???? what job do you do?

OpenStudy (anonymous):

psh, i wish they paid be like they loved and needed me lolol. Im trying to get a position at the college instead. so i can go to class and work in the same place

OpenStudy (akshay_budhkar):

Small college..lol

OpenStudy (akshay_budhkar):

all the best.. you can be online on your mobile though

OpenStudy (anonymous):

Sounds like the high school I work at. or used to work at...i guess

OpenStudy (akshay_budhkar):

USED TO?

OpenStudy (anonymous):

its a "low-tier" high school, where over half the students struggle to pass standardized testing.

OpenStudy (anonymous):

i want to put in my two weeks <.< lol i want to tutor at the college instead.

OpenStudy (akshay_budhkar):

Are you the reason for such results Joe? lol!!

OpenStudy (anonymous):

im the reason for good results! i work with kids with learning disabilities. i make them pro! lol

OpenStudy (akshay_budhkar):

lol obedient they are!

OpenStudy (anonymous):

i dont tell me students about this site <.< i spend too much time here.....this is my away place lol

OpenStudy (anonymous):

maybe i should change my pic....and my screen name.....then it would be ok lol

OpenStudy (akshay_budhkar):

LOL

OpenStudy (akshay_budhkar):

You guys like teaching?

OpenStudy (akshay_budhkar):

then why do you do it?

OpenStudy (anonymous):

i enjoy the act of teaching, i dont enjoy working as a teacher.

OpenStudy (anonymous):

for a job? gotta pay those bills. There arent millions of jobs in the economy anymore, gotta take what you can get.

OpenStudy (akshay_budhkar):

that is confusing joe what you said?

OpenStudy (akshay_budhkar):

oh Right

OpenStudy (anonymous):

Like, what we do here, i enjoy, we are helping people, answering questions, and learning. Even if we arent getting paid, i enjoy this.

OpenStudy (akshay_budhkar):

national science academy?

OpenStudy (anonymous):

But working at a high school, theres too much regulation and rules, meetings, requirements. The administration is horrible. etc etc. its a drain to work.

OpenStudy (akshay_budhkar):

oh cool

OpenStudy (anonymous):

thats right myininaya

OpenStudy (akshay_budhkar):

you guys are great.. hats off!

OpenStudy (anonymous):

i want to be locked in a basement room. just number crunching. every couple of hours a small door will open and someone will toss in a 3 day old blueberry muffin.

OpenStudy (anonymous):

dream job lol

OpenStudy (akshay_budhkar):

LOL Cool job

OpenStudy (akshay_budhkar):

myininaya you from US too?

OpenStudy (akshay_budhkar):

cool! all the best for the meeting

OpenStudy (akshay_budhkar):

lol

myininaya (myininaya):

there i deleted everything i said so if anyone would had figured out my identity they can't now

OpenStudy (anonymous):

lol, make it look like we were talking to ourselves!

myininaya (myininaya):

lol

OpenStudy (anonymous):

I'm stumped! thank you!!!!

myininaya (myininaya):

\[\frac{\pi}{2}-(2x+45)=30-x \] \[\frac{\pi}{2}-45-30=-x+2x\] \[x=\frac{\pi}{2}-75\] but your question says use the factor theorem ...

myininaya (myininaya):

and somehow we need to find the minimal value for x